# The 18 May 2024 superbolide over the Iberian Peninsula: USG space sensors and ground-based independent observations

### v1

@@ Abstract
On 18 May 2024, a superbolide traversed the western part of the Iberian Peninsula, culminating its flight over the Atlantic Ocean and generating significant media attention. This event was caused by a weak carbonaceous meteoroid of 93.0$\pm$0.8 cm, with a density of 1613$\pm$12 kg m-3, entering the atmosphere at 40.1$\pm$0.4 km s-1 with an angle of 10.93$\pm$0.02. The luminous phase started at 137.88$\pm$0.05 km and ended at an altitude of 53.78$\pm$0.07 km. The meteoroid’s heliocentric orbit was characterized by an inclination of 14.42$\pm$0.23, a high eccentricity of 0.950$\pm$0.004, a semi-major axis of 2.31$\pm$0.13 au, and a notably short perihelion distance of 0.116$\pm$0.003 au. The superbolide was recorded by multiple ground-based stations of the Spanish Meteor Network (SPMN), the European Space Agency (ESA), and the U.S. Government (USG) space sensors. Our analysis shows a good agreement with the radiant and velocity data reported by the Center for Near-Earth Object Studies (CNEOS), with a deviation of 1.6 and -0.4 km s-1, respectively. Due to the absence of observable deceleration, we successfully reconciled satellite radiometric data with a purely dynamic atmospheric flight model, constraining the meteoroid’s mass and coherently fitting its velocity profile.
@@ 2.1 Ground-based observation
We use our 3D-FireTOC software to analyze the event. This tool integrates computer vision and machine learning for detecting and tracking fireballs, robust camera calibration to correct distortions, the plane intersection method for triangulation, and heliocentric orbit determination (Peña-Asensio et al., 2021b, a, 2023b, 2023a). Using a model of atmospheric mass density, the dynamic strength where the meteoroid disruption occurs is computed as $S=\rho_{atm}v^{2}$ (Bronshten, 1983). From that, thanks to the empirical strength-density relation established by Collins et al. (2005), we calculate the impactor density as:
@@
At the onset of the luminous phase and down to altitudes near 60 km, the atmosphere is sufficiently thin that deceleration is practically negligible. Consequently, all observed points during this phase equally represent the initial velocity, allowing for a robust estimate.
@@ 2.2 Space-based observation
As the superbolide did not penetrate deeply into the atmosphere and was observed from long distances, the ground-based measurements are insufficient for fitting a dynamic deceleration model. Therefore, we use as a proxy the total radiated energy ($E$) measured by the USG sensors, which has proven reliable. By assuming that the energy recorded is equal to the kinetic energy, we estimate the mass of the projectile as:
$m_{0}=\frac{2\cdot E}{v^{2}}.$ (2)
@@
With $\alpha$ determined, we fit the atmospheric flight model to the observed data to obtain $\beta$ (Gritsevich, 2007):
@@ 3 Results and comparison
Initial longitude [] $\lambda_{0}$ 5.5197$\pm$0.0011 W
Initial latitude [] $\varphi_{0}$ 38.5777$\pm$0.0012 N
Initial velocity [km s-1] $v_{0}$ 40.1$\pm$0.4
@@
Energy peak longitude [] $\lambda_{p}$ 8.2930$\pm$0.0014 W 8.8 W -0.507
Energy peak latitude [] $\varphi_{p}$ 40.8852$\pm$0.0011 N 41.0 N -0.1148
Energy peak velocity [km s-1] $v_{p}$ 40.1$\pm$0.1 40.4 -0.3
@@
Dynamic strength [kPa] $S$ 41.0$\pm$0.9 72.41 -31.41
Meteoroid density [kg m-3] $\rho_{m}$ 1613$\pm$12 1946 -333
Meteoroid mass [kg] $m_{0}$ 678$\pm$13 666 12
Initial diameter [cm] $D$ 93.0$\pm$0.8 87 6
Geo. velocity [km s-1] $v_{R}$ 38.2$\pm$0.4 38.6 -0.4
Geo. radiant R.A. [] $\alpha_{R}$ 261.65$\pm$0.05 262.4 -0.75
Geo. radiant Dec. [] $\delta_{R}$28.26$\pm$0.06 -29.7 1.44
Semi-major axis [au] $a$ 2.31$\pm$0.13 2.3 0.01
Eccentricity $e$ 0.950$\pm$0.004 0.95 0
Inclination [] $i$ 14.42$\pm$0.23 18.2 -3.78
Argument of periapsis [] $\omega$ 144.87$\pm$0.23 145.4 -0.53
Long. of the asc. node [] $\Omega$ 238.1253$\pm$0.0001 238.1321 -0.0068
Perihelion distance [au] $q$ 0.116$\pm$0.003 0.11 0.006
True anomaly [] $f$ 215.13$\pm$0.24 214.6 0.53
Period [year] $P$ 3.51$\pm$0.29 3.5 0.01
Tisserand’s parameter $Tj$ 2.66$\pm$0.13 2.7 -0.04
Ballistic coefficient $\alpha$ 17.07$\pm$0.06
Mass-loss parameter $\beta$ 0.95$\pm$0.10
@@
The search for possible associations with meteoroid streams or parent bodies for superbolide SPMN180524F yielded negative results using both traditional D-criteria and Machine Learning distance metrics (Peña-Asensio & Sánchez-Lozano, 2024). If the SPMN180524F superbolide resulted from a cometary disruption, additional fragments could potentially reach the Earth. Nevertheless, these are difficult to detect in advance due to the typical low albedo of carbonaceous chondrites (Trigo-Rodríguez et al., 2014; Tanbakouei et al., 2020). Given the estimated meteoroid size of approximately 0.93 m and a density of 1613 kg m-3, we hypothesize its origin might be the disruption of a weak carbonaceous body. Remarkably, a recent study of the samples returned by the OSIRIS-REx NASA mission revealed similar bulk density values (Lauretta et al., 2024), which could suggest a Bennu-like parent body.
@@
The scenario of SPMN180524F originating from a disrupted comet is plausible, as its orbital eccentricity pulled our meteoroid to the outer asteroid belt, not far from Jupiter. The plausible connection between carbonaceous chondrites (CCs) and comets, long suggested, is further supported by recent studies of the reflectance spectra of comet 2P/Encke and ungrouped CCs (Tanbakouei et al., 2020). The role of erosion and subsequent dehydration caused by thermal processing of cometary nuclei was elucidated by Rosetta’s study of 67P/Churyumov-Gerasimenko, providing insights into the evolution of meteoroids resulting from their disruption (Fulle et al., 2020; Koschny et al., 2019; Trigo-Rodríguez et al., 2019). Researches on cometary formation and disintegration products suggest that centimeter-sized pebbles with higher density can be preserved in the interior of comets (Schräpler et al., 2022; Trigo-Rodríguez & Blum, 2022). These pebbles can survive at higher dynamic strengths than the fine-grained materials forming the meteoroid interstitial matrix.
@@
There are examples of superbolides produced by carbonaceous chondrite (CC) projectiles, such as the Tagish Lake meteorite fall, which had an estimated meteoroid density of 1500 kg m-3 but resulted in higher density ungrouped CCs (Brown et al., 2000), as well as the Maribo, Sutter’s Mill, Flensburg, and Winchcombe meteorites (Haack et al., 2012; Jenniskens et al., 2012; Borovička et al., 2021; McMullan et al., 2024). In any case, these fragile meteoroids often generate extremely bright bolides without yielding noticeable meteorites. Over more than 25 years of operation, our SPMN network has recorded several cometary superbolides; one of the most memorable events occurred on July 11, 2008, which, due to its anomalous orbit, was tentatively associated with the disruption of comet C/1919 Q2 Metcalf. (Trigo-Rodríguez et al., 2009).
@@
The superbolide SPMN180524F belongs to the group of events well-measured from space, as the geocentric radiant is deviated in 1.6 and the geocentric velocity in just -0.4 km s-1, resulting in a good agreement on the heliocentric orbit. One notable divergence is the 4.43 of disagreement in slope, which yields a -3.78 discrepancy in the orbital inclination. It is worth noting the difference of -333 kg m-3 in density, mainly due to the 3.61 km of disagreement in the height where the radiation peak occurs.
@@ 4 Conclusions
The 18 May 2024 superbolide was a unique event that showcased how a meter-sized meteoroid can produce a spectacular display of color and luminosity. It exemplified nicely how the Earth’s atmosphere is an excellent shield for meter-sized projectiles. A detailed analysis of the superbolide characteristics suggests that the meteoroid likely originated from the disruption of a weak carbonaceous body or a transitional object, potentially explaining its eccentric orbit with aphelia in the outer asteroid belt. Its short perihelion distance suggests a recent detachment from its parent body, as the long-term survival of fragile materials would be challenging under such thermal stress conditions close to the Sun. Given the high velocity and fragmentation observed, meteorite survival is unlikely. However, the previous Maribo carbonaceous chondrite fall demonstrated that even weak meteoroids can produce small meteorites at relatively high speeds (Borovička et al., 2019).
Our analysis demonstrated strong concordance with the data reported by CNEOS both in radiant and velocity, and subsequently in the heliocentric orbital elements, with the speed and azimuth being very well measured, but a non-negligible discrepancy in the slope. We reconciled the satellite radiometric data with a purely dynamic atmospheric flight model to constrain the meteoroid’s mass and consistently derive the atmospheric deceleration curve.
From an impact risk perspective, the event raises questions about why a meter-sized meteoroid was not detected by current telescopic surveys, especially considering that these minor bodies monitoring programs have successfully identified some asteroids of just a few meters in diameter before their collision with Earth. Our results provide a clear explanation: the meteoroid was simply too small and exhibited a low albedo to be detectable.

### v2

@@ Abstract
On 18 May 2024, a superbolide traversed the western part of the Iberian Peninsula, culminating its flight over the Atlantic Ocean and generating significant media attention. This event was caused by a weak carbonaceous meteoroid of 89.8$\pm$0.4 cm, with a density of 1660 kg m-3, entering the atmosphere at 41.6 km s-1 with an angle of 10.9. The luminous phase started at 138 km and ended at an altitude of 54 km. The meteoroid’s heliocentric orbit was characterized by an inclination of 15.4, a high eccentricity of 0.965, a semi-major axis of 3 au, and a notably short perihelion distance of 0.11 au. The superbolide was recorded by multiple ground-based stations of the Spanish Meteor Network (SPMN), the European Space Agency (ESA), and the U.S. Government (USG) space sensors. Our analysis shows a relatively good agreement with the radiant and velocity data reported by the Center for Near-Earth Object Studies (CNEOS), with a deviation of 1 and 1.2 km s-1, respectively. Due to the absence of observable deceleration, we successfully reconciled satellite radiometric data with a purely dynamic atmospheric flight model, constraining the meteoroid’s mass and coherently fitting its velocity profile. The inferred bulk density and aerodynamic strength of the meteoroid from the flight model are compatible with the properties recently measured for sample-returned materials from asteroid Bennu. The physical properties and the orbital results suggest that this meteoroid originated from a recent disruption of a comet, indicating the existence of hazardous meter-sized projectiles arriving on Earth from objects formed in the distant regions of the Solar System.
@@ 2.1 Ground-based observation
We use our 3D-FireTOC software to analyze the event. This tool integrates computer vision and machine learning for detecting and tracking fireballs, robust camera calibration to correct distortions, the plane intersection method for triangulation, and heliocentric orbit determination (Peña-Asensio et al., 2021b, a, 2023b, 2023a). Using a model of atmospheric mass density and the measured velocity, the dynamic strength where the meteoroid disruption occurs is computed as $S=\rho_{atm}v^{2}$ (Bronshten, 1983). From that, thanks to the empirical strength-density relation established by Collins et al. (2005), we calculate the impactor density as:
@@
At the onset of the luminous phase and down to altitudes near 60 km, the atmosphere is sufficiently thin that deceleration is barely observable. Consequently, all measured points during this phase equally represent the initial velocity, allowing for a robust estimate.
@@ 2.2 Space-based observation
As the superbolide did not penetrate deeply into the atmosphere and was observed from long distances, the ground-based measurements are insufficient for fitting a dynamic deceleration model. Therefore, we use as a proxy the total radiated energy ($E$) measured by the USG sensors, which has proven reliable. By assuming that the energy recorded is equal to the kinetic energy, we estimate the initial mass of the projectile as:
$m_{0}=\frac{2E}{v^{2}}.$ (2)
@@
By assuming the terminal mass is equal to zero, $\beta$ can be estimated from the terminal height $h_{e}$ and $\alpha$ (Moreno-Ibáñez et al., 2015):
$\beta=\frac{e^{h_{e}/h_{0}}}{2\alpha}.$ (4)
With $\alpha$ and $\beta$ determined, we fit the atmospheric flight model to the observed data to obtain the velocity profile (Gritsevich, 2007):
@@ 3 Results and comparison
Initial longitude [] $\lambda_{0}$ 5.5197$\pm$0.0010 W
Initial latitude [] $\varphi_{0}$ 38.5777$\pm$0.0011 N
Initial velocity [km s-1] $v_{0}$ 41.64$\pm$0.19
@@
Energy peak longitude [] $\lambda_{p}$ 8.2932$\pm$0.0017 W 8.8 W -0.5068
Energy peak latitude [] $\varphi_{p}$ 40.8851$\pm$0.0015 N 41.0 N -0.1149
Energy peak velocity [km s-1] $v_{p}$ 41.63$\pm$0.19 40.4 1.23
@@
Dynamic strength [kPa] $S$ 44.4$\pm$0.4 72.41 -28.01
Meteoroid density [kg m-3] $\rho_{m}$ 1658$\pm$6 1946 -288
Meteoroid mass [kg] $m_{0}$ 628$\pm$6 666 -38
Initial diameter [cm] $D$ 89.8$\pm$0.4 87 2.8
Geo. velocity [km s-1] $v_{R}$ 39.87$\pm$0.19 38.6 1.27
Geo. radiant R.A. [] $\alpha_{R}$ 261.56$\pm$0.04 262.4 -0.84
Geo. radiant Dec. [] $\delta_{R}$28.13$\pm$0.04 -29.7 1.57
Semi-major axis [au] $a$ 3.01$\pm$0.11 2.3 0.71
Eccentricity $e$ 0.965$\pm$0.002 0.95 0.015
Inclination [] $i$ 15.39$\pm$0.12 18.2 -2.81
Argument of periapsis [] $\omega$ 145.52$\pm$0.12 145.4 0.12
Long. of the asc. node [] $\Omega$ 238.1257$\pm$0.0001 238.1321 -0.0064
Perihelion distance [au] $q$ 0.106$\pm$0.001 0.11 -0.004
True anomaly [] $f$ 214.49$\pm$0.12 214.6 -0.11
Period [year] $P$ 5.2$\pm$0.3 3.5 1.7
Tisserand’s parameter $Tj$ 2.1$\pm$0.1 2.7 -0.6
Ballistic coefficient $\alpha$ 17.22$\pm$0.05
Mass-loss parameter $\beta$ 53.1$\pm$0.4
@@
The search for possible associations with meteoroid streams or parent bodies for superbolide SPMN180524F yielded negative results using both traditional D-criteria and Machine Learning distance metrics (Peña-Asensio & Sánchez-Lozano, 2024). If the SPMN180524F superbolide resulted from a cometary disruption, additional fragments could potentially reach the Earth. Nevertheless, these are difficult to detect in advance due to the typical low albedo of carbonaceous chondrites (Trigo-Rodríguez et al., 2014; Tanbakouei et al., 2020). Given the estimated meteoroid size of approximately 0.9 m and a density of 1660 kg m-3, we hypothesize its origin might be the disruption of a comet. Remarkably, a recent study of the samples returned by the OSIRIS-REx NASA mission revealed similar bulk density values (Lauretta et al., 2024), which could suggest a Bennu-like parent body.
@@
The scenario of SPMN180524F originating from a disrupted comet is plausible, as its orbital eccentricity carried the meteoroid well beyond the orbit of Jupiter. The plausible connection between carbonaceous chondrites (CCs) and comets, long suggested, is further supported by recent studies of the reflectance spectra of comet 2P/Encke and ungrouped CCs (Tanbakouei et al., 2020). The role of erosion and subsequent dehydration caused by thermal processing of cometary nuclei was elucidated by Rosetta’s study of 67P/Churyumov-Gerasimenko, providing insights into the evolution of meteoroids resulting from their disruption (Fulle et al., 2020; Koschny et al., 2019; Trigo-Rodríguez et al., 2019). Researches on cometary formation and disintegration products suggest that centimeter-sized pebbles with higher density can be preserved in the interior of comets (Schräpler et al., 2022; Trigo-Rodríguez & Blum, 2022). These pebbles can survive at higher dynamic strengths than the fine-grained materials forming the meteoroid interstitial matrix.
@@
There are examples of superbolides produced by carbonaceous chondrite (CC) projectiles, such as the Tagish Lake meteorite fall, which had an estimated meteoroid density of 1500 kg m-3 but resulted in higher density ungrouped CCs (Brown et al., 2000), as well as the Maribo, Sutter’s Mill, Flensburg, and Winchcombe meteorites (Haack et al., 2012; Jenniskens et al., 2012; Borovička et al., 2021; McMullan et al., 2024). In any case, these fragile meteoroids often generate extremely bright bolides without yielding noticeable meteorites (Borovička et al., 2022). Over more than 25 years of operation, our SPMN network has recorded several cometary superbolides; one of the most memorable events occurred on July 11, 2008, which, due to its anomalous orbit, was tentatively associated with the disruption of comet C/1919 Q2 Metcalf. (Trigo-Rodríguez et al., 2009).
@@
The superbolide SPMN180524F belongs to the group of events well-measured from space, as the apparent radiant is deviated in 0.98 and the velocity at the energy peak in 1.23 km s-1, resulting in a relatively good agreement on the heliocentric orbit. This difference in velocity leads to a 0.71 au dissent in the semi-major axis. One notable divergence is the 4.43 of disagreement in slope, which yields a -2.81 discrepancy in the orbital inclination. It is worth noting the difference of -288 kg m-3 in density, mainly due to the 3.61 km of disagreement in the height where the radiation peak occurs.
@@ 4 Conclusions
The 18 May 2024 superbolide was a unique event that showcased how a fragile meter-sized meteoroid can produce a spectacular display of color and luminosity, and it exemplified how the Earth’s atmosphere is an excellent shield for this type of impactor. A detailed analysis of the superbolide characteristics suggests that the meteoroid likely originated from the disruption of an eccentric comet, potentially explaining its relatively unstable orbit crossing the orbit of Jupiter. In any case, its short perihelion distance also suggests a recent detachment from its parent comet, as the long-term survival of fragile materials would be challenging under such thermal stress conditions close to the Sun. Given the high velocity and fragmentation observed, a meteorite’s survival seems unlikely. However, the previous Maribo carbonaceous chondrite fall demonstrated that even weak meteoroids can produce small meteorites at relatively high speeds (Borovička et al., 2019).
Our analysis demonstrated relatively good agreement with the data reported by CNEOS both in radiant and velocity, and subsequently in the heliocentric orbital elements. We reconciled the satellite radiometric data with a purely dynamic atmospheric flight model to constrain the meteoroid’s mass and consistently derive the atmospheric deceleration curve.
From an impact risk perspective, the event raises questions about why a meter-sized meteoroid was not detected by current telescopic surveys, especially considering that these small object monitoring programs have successfully identified some asteroids of just a few meters in diameter before their collision with Earth. Our results provide a clear explanation: the meteoroid was simply too small and exhibited a low albedo to be detectable.